### v1

# thanks: These authors contribute equally to this workthanks: These authors contribute equally to this workthanks: These authors contribute equally to this work

Abstract
The emergence of the arrow of time in quantum many-body systems stems from the inherent tendency of Hamiltonian evolution to scramble quantum information and increase entanglement. While, in principle, one might counteract this temporal directionality by engineering a perfectly inverted Hamiltonian to reverse entanglement growth, such a scenario is fundamentally unstable because even minor imperfections in the backward evolution can be exponentially amplified, a hallmark of quantum many-body chaos. Therefore, successfully reversing quantum many-body dynamics demands a deep understanding of the underlying structure of quantum information scrambling and chaotic dynamics. Here, by using solid-state nuclear magnetic resonance on a macroscopic ensemble of randomly interacting spins, we measure the out-of-time-ordered correlator (OTOC) and validate key predictions of scramblon theory, a universal theoretical framework for information scrambling. Crucially, this theory enables us to isolate and mitigate errors in the OTOC caused by imperfections in the backward evolution. As a result, this protocol uncovers the anticipated exponential behavior of quantum many-body chaos and extracts the quantum Lyapunov exponent in a many-body experimental system for the first time. Our results push the fundamental limits of dynamical reversibility of complex quantum systems, with implications for quantum simulation and metrology.
Quantum information scrambling lies at the heart of thermalization in isolated quantum systems [1, 2, 3, 4, 5]. When a low-entanglement initial state evolves under a Hamiltonian $\hat{H}$, local information spreads across the entire system, becoming irretrievable via local measurements. At the same time, the entanglement entropy grows, rendering the system locally a thermal state. Intriguingly, one can envision a gedanken experiment: if we apply the reversed Hamiltonian $-\hat{H}$ to this thermal state, is it possible to recover the original low-entanglement state and its local information? This question echoes a profound analogy in black hole physics: whether information thrown in and thermalized by a black hole can ever be retrieved. Now, advances in quantum control techniques on various synthetic quantum platforms attempt to turn such a thought experiment into a tangible ambition [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], raising an even deeper question: can engineered inverted dynamics defy the arrow of time?
In realistic quantum systems, however, such an oversimplified scenario does not materialize. As illustrated by Fig. 1a, even infinitesimal imperfections in the engineered reversed Hamiltonian can lead to an exponentially amplified error as evolution time increases, a hallmark of quantum many-body chaos [32, 33, 34, 35, 36]. Given that no practical implementation can perfectly realize $-\hat{H}$, the inherent chaotic nature of quantum many-body dynamics fundamentally prohibits the retrieval of information from a thermal state. The same reasoning extends analogously to the challenge of recovering information from a black hole. Recent experimental observations, spanning nuclear magnetic resonance (NMR) systems [6, 7, 8, 9, 10, 11, 12, 13, 14], trapped ions [15, 16, 17, 18], superconducting qubits [19, 20, 21, 22, 23], ultracold atoms [24, 25, 26, 27, 28, 29, 30], and nitrogen vacancy centers [31], have also directly observed the inevitable effects of imperfect backward evolution.
Figure 1: Schematic of Experimental Protocol. (a) Chaotic quantum many-body dynamics: Unitary evolution under $\hat{H}$ scrambles quantum information, transforming a low-entanglement initial state into a highly entangled state. Under time-reversed dynamics with perturbed Hamiltonian $-\hat{H}+\delta\hat{H}$, the system exhibits exponential deviation from perfect state recovery. (b) Microscopic structure of adamantane (C10H16) powder: Each granule contains adamantane molecules arranged in a face-centered cubic lattice. Individual molecules consist of spin-$1/2$ 1H nuclei and spinless 12C atoms. (c) Top: Floquet pulse sequence converting the dipolar Hamiltonian Eq. (1) into the engineered form Eq. (2). Bottom: Experimental sequence for MQC and OTOC measurements.
Although errors in time-reversed dynamics are inherently unavoidable, a potential solution is to post-process experimental data to extrapolate an error-free limit. However, achieving an unbiased extrapolation remains highly challenging unless the imperfections meet the following two criteria: (1) the error model in the backward evolution must follow a universal behavior, independent of microscopic details; and (2) this universal behavior can be captured by a theoretical framework that yields an analytical ansatz. Such an ansatz would enable both the reliable quantification of imperfections in the backward evolution and robust extrapolation to the error-free limit.
Remarkably, these two conditions can be simultaneously satisfied across a broad class of systems. Quantum information scrambling and many-body chaos are intimately linked, both characterized by the OTOC [37, 32, 33, 34, 36, 35]. The recently developed scramblon theory has identified the scramblon, a novel type of collective excitation in quantum many-body systems with all-to-all connectivity, as the fundamental carrier of scrambling dynamics [38, 39, 40, 41]. Analogous to phonons mediating density fluctuations or spin waves governing spin dynamics, scramblons dictate the spread of quantum information. The scramblon propagator grows exponentially with time, outpacing all other modes and dominating the late-time behavior of the OTOC, thereby resulting in fast information scrambling and a universal exponential behavior of the OTOC. Notably, in systems with holographic duality, scramblons in the boundary quantum theory correspond to gravitons in the bulk gravity [38, 42].
In this work, we measure the OTOC in a solid-state NMR sample with a macroscopic number of randomly interacting spins, a system beyond the capabilities of classical simulation. This can be achieved using a well-established protocol for measuring the multiple quantum coherence (MQC) spectrum (Fig. 1c, bottom panel) [43, 44]. The measurement involves the reverse evolution of this many-body system, which, of course, is imperfect. We utilize scramblon theory to derive a fitting ansatz for the measured OTOC, achieving good agreement with the experiment. Moreover, the scramblon theory imposes several strong constraints on the fitting parameters, and we verify that these constraints are well satisfied by the fitting results. In other words, the experimental data strongly validate the scramblon theory. With this fitting formula, we can extrapolate the measured OTOC to the error-free limit. We demonstrate that the OTOC after this error mitigation exhibits the anticipated long-time behavior, in sharp contrast to other methods explored in the previous literature.
Figure 2: Fitting Experimental Data with the Scramblon Ansatz. Three typical experimental data $F^{(\lambda)}_{\alpha\gamma}(\phi,t)$ for three different $\phi=0$, $\pi/64$ and $\pi/32$. The superscript $\lambda$ labels three different Hamiltonians, as described in the main text. The solid curves are fitting results obtained using the scramblon ansatz. Experimentally, the $95\%$ confidence intervals determined from read-out noise are approximately $10^{-4}$, and therefore the error bars are contained within the data markers.
The experiment is performed on a powdered adamantane (C10H16) sample, as illustrated in Fig. 1b and detailed in the supplementary material [45]. Each adamantane molecule contains 16 nuclear spin-$1/2$ hydrogen nuclei, denoted by $\mathbf{S}_{im}$, where the index $i$ labels the molecules and $m=1,\dots,16$ identifies individual spins within a molecule. Due to rapid molecular thermal motion, intramolecular spin interactions are averaged out, leaving only the intermolecular dipolar coupling as the dominant spin-spin interaction. The sample is placed in a uniform $9.4$ T magnetic field, which aligns the spin quantization axes and induces fast Larmor precession of all spins. In the rotating frame and under the secular approximation, the effective dipolar Hamiltonian takes the form [46]:
$\hat{H}_{\text{d}}=\sum_{i<j,m,n}J_{ij}\left(-\hat{S}^{x}_{im}\hat{S}^{x}_{jn}-\hat{S}^{y}_{im}\hat{S}^{y}_{jn}+2\hat{S}^{z}_{im}\hat{S}^{z}_{jn}\right).$ (1)
Here, the coupling strengths $J_{ij}$ are treated as random variables with zero mean ($\overline{J_{ij}}=0$) and variance $\overline{J^{2}_{ij}}=4J^{2}/N$, where $N$ is the total number of spins and $J\sim 2\pi\times 1460$ Hz [14]. This randomness of $J_{ij}$ arises from the anisotropic nature of dipolar interaction and the random crystalline orientations in the powder sample [45].
Then, by applying periodic global radio-frequency pulse sequences to rotating spins, as depicted in Fig. 1c, a more general effective Hamiltonian can be engineered via the Floquet-Magnus expansion as $\hat{H}_{\text{eff}}=\hat{H}_{0}+\hat{H}_{1}$. $\hat{H}_{0}$ represents the leading-order contribution in the expansion, which can be generally written as:
$\hat{H}_{0}=\sum_{i<j,m,n}\sum_{\mu,\nu}J_{ij}\xi_{\mu\nu}\hat{S}^{\mu}_{im}\hat{S}^{\nu}_{jn},$ (2)
where $\mu,\nu=x,y,z$. The protocol generating Eq. (2) imposes specific constraints on the coefficients $\xi_{\mu\nu}$, namely the traceless condition $\sum_{\mu=x,y,z}\xi_{\mu\mu}=0$ and the symmetry condition $\xi_{\mu\nu}=\xi_{\nu\mu}$. An important observation is that if a set $\{\xi_{\mu\nu}\}$ satisfies these constraints, so does the set $\{-\xi_{\mu\nu}\}$. This implies that Floquet engineering always enables reversal of $\hat{H}_{0}$. However, the higher-order terms, denoted by $\hat{H}_{1}$, generally cannot be simultaneously reversed. Therefore, without loss of generality, we denote the forward evolution as governed by $\hat{H}=\hat{H}_{0}+\hat{H}_{1}$, and the backward evolution as governed by $-\hat{H}_{0}-\hat{H}_{1}^{\prime}=-\hat{H}-\delta\hat{H}$. Here, $\delta\hat{H}$, representing the difference between $\hat{H}_{1}$ and $\hat{H}^{\prime}_{1}$, is the imperfection in the reversed dynamics. In this system, $\hat{H}_{1}$ is typically on the order of a few percent relative to $\hat{H}_{0}$ [14].
The measurement protocol is schematically illustrated in Fig. 1(c). We begin by defining the total spin operator along the $\gamma$-axis as $\hat{O}_{\gamma}=\sum_{im}\hat{S}^{\gamma}_{im}$, where $\gamma\in\{x,y,z\}$. The system is initialized in a weakly polarized state described by the density matrix $\hat{\rho}_{\gamma}\propto\mathbbm{1}+\epsilon\hat{O}_{\gamma}$, with $\epsilon\ll 1$ characterizing the small polarization. Following forward time evolution under the Hamiltonian $\hat{H}$, the density matrix transforms as $\hat{\rho}_{\gamma}(t)={\rm e}^{-{\rm i}\hat{H}t}\hat{\rho}_{\gamma}{\rm e}^{{\rm i}\hat{H}t}$. Subsequently, we implement a uniform spin rotation by an angle $\phi$ around the $\alpha$-axis ($\alpha\in\{x,y,z\}$), yielding the transformed density matrix: $\hat{\rho}_{\alpha\gamma}(\phi,t)={\rm e}^{-{\rm i}\hat{O}_{\alpha}\phi}{\rm e}^{-{\rm i}\hat{H}t}\hat{\rho}_{\gamma}{\rm e}^{{\rm i}\hat{H}t}{\rm e}^{{\rm i}\hat{O}_{\alpha}\phi}$. This is followed by backward evolution as described above. Finally, we measure the total magnetization along the $\gamma$-direction $\langle\hat{O}_{\gamma}\rangle$, which equals the correlation function $F_{\alpha\gamma}(\phi,t)$ defined as:
$F_{\alpha\gamma}(\phi,t)=\text{Tr}[\hat{O}_{\gamma}(0)\hat{\mathcal{V}}^{\dagger}(t){\rm e}^{-{\rm i}\phi\hat{O}_{\alpha}(t)}\hat{O}_{\gamma}(0){\rm e}^{{\rm i}\phi\hat{O}_{\alpha}(t)}\hat{\mathcal{V}}(t)].$ (3)
Here, $\hat{\mathcal{V}}(t)\equiv\exp\left(-{\rm i}\mathbf{T}\int_{0}^{t}{\rm d}t^{\prime}\delta\hat{H}(t^{\prime})\right)$ represents the time-evolution operator in the interaction picture, with $\mathbf{T}$ denoting time ordering. All operators $\hat{O}(t)$ and $\delta\hat{H}(t)$ evolve according to the Heisenberg picture under $\hat{H}$. For normalization, we divide $F_{\alpha\gamma}(\phi,t)$ by $\mathcal{C}=\langle\hat{O}_{\gamma}^{2}\rangle=N/4$ to ensure $F_{\alpha\gamma}(0,0)=1$. The MQC spectrum can be obtained through the Fourier transformation of $F_{\alpha\gamma}(\phi,t)$ with respect to $\phi$ [43, 44].
We now highlight the relationship between $F_{\alpha\gamma}(\phi,t)$ and OTOCs through two limits: (i) For perfect reversed dynamics ($\delta\hat{H}=0$), $F_{\alpha\gamma}(\phi,t)$ reduces to
$F_{\alpha\gamma}^{(0)}(\phi,t)=\text{Tr}[\hat{O}_{\gamma}(0){\rm e}^{-{\rm i}\phi\hat{O}_{\alpha}(t)}\hat{O}_{\gamma}(0){\rm e}^{{\rm i}\phi\hat{O}_{\alpha}(t)}],$ (4)
representing an OTOC of operators ${\rm e}^{{\rm i}\phi\hat{O}_{\alpha}(t)}$ and $\hat{O}_{\gamma}$; (ii) At $\phi=0$, it becomes
$F_{\alpha\gamma}(0,t)=\text{Tr}[\hat{O}_{\gamma}(0)\hat{\mathcal{V}}^{\dagger}(t)\hat{O}_{\gamma}(0)\hat{\mathcal{V}}(t)],$ (5)
and this OTOC, also known as the Loschmidt echo, probes imperfections in the reversed dynamics. Both cases can be computed by scramblon theory [41], strongly suggesting that the general behavior of $F_{\alpha\gamma}(\phi,t)$, although more challenging to derive, can also be captured by scramblon theory. Indeed, as we show in the Supplementary Material [45], an ansatz for $F_{\alpha\gamma}(\phi,t)$ in the small-$\phi$ regime can be derived as:
$F_{\alpha\gamma}(\phi,t)=\frac{1}{\left(1+a{\rm e}^{\varkappa t}+b\phi^{2}{\rm e}^{\varkappa t}\right)^{2\Delta}},$ (6)
where $\{a,b,\varkappa,\Delta\}$ are all fitting parameters. The $\phi=0$ limit of Eq. (6) reproduces the Loschmidt echo (Eq. (5)), with $a\neq 0$ quantifying the reversal imperfection. We note that a recent theoretical work also discussed a similar issue in a specific solvable model [47].
Figure 3: Verifying the Predictions of Scramblon Theory. The three columns correspond to the fitting parameters extracted from three different Hamiltonians: $\hat{H}^{(1)}$, $\hat{H}^{(2)}$, and $\hat{H}^{(3)}$. Subfigures a-c in the upper row display the fitting parameters for different values of $\alpha$ and $\gamma$, with green diamonds representing $a\times 50$, yellow squares representing $\Delta$, and blue triangles representing $\varkappa$. The dashed lines represent $\varkappa$ averaged over $\alpha$ and $\gamma$, and $a$ and $\Delta$ averaged over $\alpha$. Subfigures d-f in the lower row show the eigenvalues $\sigma_{1},\sigma_{2},\sigma_{3}$ for the symmetric matrix $M^{\text{s}}$ and the absolute value of eigenvalues $\sigma_{4},\sigma_{5},\sigma_{6}$ for the antisymmetric matrix $M^{\text{a}}$. The error bars on the fitting parameters represent the $95\%$ confidence interval, including the standard deviation from each fitting residual $\sigma_{\text{res}}$ and the uncertainty arising from the choice of fitting range $\sigma_{\text{range}}$ [45].
We investigate three representative chaotic Hamiltonians: (i) $\hat{H}_{0}^{(1)}$ with $\xi_{xx}=-\xi_{yy}=0.4$, (ii) $\hat{H}_{0}^{(2)}$ with $\xi_{xy}=\xi_{yx}=0.4$, and (iii) $\hat{H}_{0}^{(3)}$ with $(\xi_{xx},\xi_{yy},\xi_{zz})=(0.025,-0.225,0.2)$, while all other $\xi_{\mu\nu}=0$. For each Hamiltonian $\hat{H}^{(\lambda)}_{0}$, we perform measurements for all nine possible combinations of $\alpha,\gamma\in\{x,y,z\}$ axes, yielding a total of $3\times 9=27$ distinct cases. For each case, we measure three different $\phi$ angles and fit the three curves jointly with a single scramblon ansatz, Eq. (6). We find that all data are excellently captured by the scramblon ansatz Eq. (6), with three representative fits shown in Fig. 2 and the complete set of all data available in the Supplementary Material [45].
The scramblon theory not only provides the fitting ansatz but also imposes strong constraints on the parameters. For a given Hamiltonian, while the fitting yields parameter sets indexed by $\alpha$ and $\gamma$, the theory predicts the following:
  1. 1.
    $\varkappa$ must be independent of both $\alpha$ and $\gamma$, since the system is dominated by a single scramblon mode and $\varkappa$ characterizes the intrinsic growth rate of this mode.
  2. 2.
    $a$ and $\Delta$ cannot depend on $\alpha$ as they are retained in the $\phi=0$ limit where no $\alpha$-axis spin rotation is applied.
  3. 3.
    The matrix $M_{\alpha\gamma}\equiv b_{\alpha\gamma}\Delta_{\gamma}$ must be symmetric and rank-1, reflecting the symmetric coupling of a single scramblon mode to both $\hat{O}_{\alpha}$ and $\hat{O}_{\gamma}$ operators.
Fig. 3 demonstrates how we verify these three constraints for our three Hamiltonians. Panels (a-c) display the values of $a$, $\Delta$, and $\varkappa$ for all nine measurement combinations per Hamiltonian, with dashed lines indicating (i) $\varkappa$ averaged over $\gamma$ and $\alpha$, and (ii) $\alpha$-averaged $a$ and $\Delta$ values that depend only on $\gamma$. Most cases show only a few percent deviations from averaged values, with a maximum deviation of $\sim 13\%$ occasionally, confirming the first two constraints. To test the third constraint, we decompose $M$ into symmetric ($M^{\text{s}}=(M+M^{\text{T}})/2$) and antisymmetric ($M^{\text{a}}=(M-M^{\text{T}})/2$) components. Eigenvalues shown in panels (d-f) reveal that $M^{\text{s}}$ exhibits only one dominant eigenvalue, and all eigenvalues of $M^{\text{a}}$ are negligible, validating the rank-1 symmetric requirement for $M$.
Figure 4: Error Mitigation of the Reversed Quantum Many-Body Dynamics. (a) $F_{\alpha\gamma}(\phi,t)$ and (b) $I_{\alpha\gamma}(t)$ defined in Eq. (8). As an example, we show $\alpha,\gamma=z,y$, $\phi=\pi/64$ and with Hamiltonian $\hat{H}^{(1)}_{0}$. The blue solid line results from the scramblon ansatz by setting $a=0$. The red triangles are error mitigated by $\tilde{F}_{\alpha\gamma}(\phi,t)$, and green circles are raw data without error mitigation. The red dash-dotted line and the green dashed line are obtained from the scramblon ansatze of $\tilde{F}_{\alpha\gamma}(\phi,t)$ and $F_{\alpha\gamma}(\phi,t)$ (without setting $a=0$). The experimental data points in (b) were extracted by fitting the corresponding data for $F_{\alpha\gamma}(\phi,t)$ with a sixth-order polynomial.
The excellent agreement with experimental data strongly validates the scramblon ansatz. Within this framework, the parameter $a$ fully captures imperfections in the time-reversed dynamics, as evidenced by $F_{\alpha\gamma}(0,t)\equiv 1$ when $a=0$. This allows an error mitigation by setting $a=0$ in Eq. (6) while keeping other fitted parameters unchanged (blue curve in Fig. 4a). For comparison, we examine an alternative mitigation approach from previous literatures that normalizes $F_{\alpha\gamma}(\phi,t)$ by $F(0,t)$ to define $\tilde{F}_{\alpha\gamma}(\phi,t)=F_{\alpha\gamma}(\phi,t)/F_{\alpha\gamma}(0,t)$ (red triangles in Fig. 4a). Within the scramblon formalism, this yields:
$\tilde{F}_{\alpha\gamma}(\phi,t)=\left(\frac{1+a{\rm e}^{\varkappa t}}{1+a{\rm e}^{\varkappa t}+b\phi^{2}{\rm e}^{\varkappa t}}\right)^{2\Delta}.$ (7)
Different approaches are compared in Fig. 4a.
To assess the reliability of different approaches, we compute $I_{\alpha\gamma}(\phi)$ from $F_{\alpha\gamma}(\phi,t)$ as:
$I_{\alpha\gamma}(t)\equiv-\partial_{\phi}^{2}F_{\alpha\gamma}(\phi,t)|_{\phi=0},$ (8)
which for perfect time reversal ($\delta\hat{H}=0$) corresponds to the OTO commutator $\langle|[\hat{O}_{\alpha}(t),\hat{O}_{\gamma}]|^{2}\rangle$ and exhibits exponential growth, reflecting the rapid increase of the operator size $\hat{O}_{\alpha}(t)$ under Heisenberg evolution before reaching the scrambling time. Analysis based on the scramblon ansatz reveals: (i) The raw data (Eq. (6)) gives $I_{\alpha\gamma}(t)=2b\Delta{\rm e}^{\varkappa t}/(1+a{\rm e}^{\varkappa t})^{1+\Delta}$, which decays at late times due to $\Delta>0$ (green dashed line in Fig. 4b); (ii) The normalized approach (Eq. (7)) yields $I_{\alpha\gamma}(t)=2b\Delta{\rm e}^{\varkappa t}/(1+a{\rm e}^{\varkappa t})$, saturating to $2b\Delta/a$ at long times (red dash-dotted line in Fig. 4b); and (iii) Only our $a=0$ mitigation produces the physically expected exponential growth $2b\Delta{\rm e}^{\varkappa t}$ (blue solid line in Fig. 4b), where $\varkappa$ emerges as the quantum Lyapunov exponent. These distinct behaviors are quantitatively compared with experimental data. To the best of our knowledge, $\varkappa$ shown in Fig. 3 serves as the first exact experimental determination of the Lyapunov exponent in a quantum many-body system with macroscopic degrees of freedom.
In conclusion, we develop an ansatz based on scramblon theory to describe reversed quantum many-body dynamics in the presence of imperfections. We validate this ansatz by fitting it to experimental data from an NMR system with macroscopically randomly interacting spins, where the stringent constraints imposed by scramblon theory are convincingly verified by the fitting outputs. This approach allows us to mitigate the effects of imperfections and recover error-resilient reversed dynamics. Notably, time-reversal-based exponential sensitivity has also been recently exploited to enhance quantum metrology in systems such as cold atoms in cavities [24, 25] and nitrogen vacancy centers [31], where imperfections also suppress this sensitivity at long evolution times. Our theory may be extended to such quantum metrology applications, potentially enabling further improvements in sensitivity.
Acknowledgements This work is supported by Innovation Program for Quantum Science and Technology 2021ZD0302005 (H.Z.), 2024ZD0300101 (P.Z.), and 2021ZD0303205 (X.P.), National Natural Science Foundation of China No. 12488301 (H.Z.), No. U23A6004 (H.Z.), No. 12374477 (P.Z.), No. 12261160569 (X.P.), the Shanghai Rising-Star Program under grant number 24QA2700300 (P.Z.), and the XPLORER Prize (H.Z. and X.P.). We thank experimental assistance from the Instruments Center for Physical Science, University of Science and Technology of China.
Data and Code Availability The raw NMR experimental data and all codes are available at [48].

### v2

@@
 Here, $\delta\hat{H}$, representing the difference between $\hat{H}_{1}$ and $\hat{H}^{\prime}_{1}$, is the imperfection in the reversed dynamics.
-In this system, $\hat{H}_{1}$ is typically on the order of a few percent relative to $\hat{H}_{0}$ [14].
+In this system, $\hat{H}_{1}$ and $\delta\hat{H}$ are typically on the order of a few percent relative to $\hat{H}_{0}$ [45, 14].
 The measurement protocol is schematically illustrated in Fig.
@@
 The error bars on the fitting parameters represent the $95\%$ confidence interval, including the standard deviation from each fitting residual $\sigma_{\text{res}}$ and the uncertainty arising from the choice of fitting range $\sigma_{\text{range}}$ [45].
-We investigate three representative chaotic Hamiltonians: (i) $\hat{H}_{0}^{(1)}$ with $\xi_{xx}=-\xi_{yy}=0.4$, (ii) $\hat{H}_{0}^{(2)}$ with $\xi_{xy}=\xi_{yx}=0.4$, and (iii) $\hat{H}_{0}^{(3)}$ with $(\xi_{xx},\xi_{yy},\xi_{zz})=(0.025,-0.225,0.2)$, while all other $\xi_{\mu\nu}=0$.
+We investigate three representative chaotic Hamiltonians: (i) $\hat{H}_{0}^{(1)}$ with $\xi_{xx}=-\xi_{yy}=0.4$, (ii) $\hat{H}_{0}^{(2)}$ with $\xi_{xy}=\xi_{yx}=0.4$, and (iii) $\hat{H}_{0}^{(3)}$ with $(\xi_{xx},\xi_{yy},\xi_{zz})=(0.2,-0.225,0.025)$, while all other $\xi_{\mu\nu}=0$.
 For each Hamiltonian $\hat{H}^{(\lambda)}_{0}$, we perform measurements for all nine possible combinations of $\alpha,\gamma\in\{x,y,z\}$ axes, yielding a total of $3\times 9=27$ distinct cases.
@@
 Panels (a-c) display the values of $a$, $\Delta$, and $\varkappa$ for all nine measurement combinations per Hamiltonian, with dashed lines indicating (i) $\varkappa$ averaged over $\gamma$ and $\alpha$, and (ii) $\alpha$-averaged $a$ and $\Delta$ values that depend only on $\gamma$.
-Most cases show only a few percent deviations from averaged values, with a maximum deviation of $\sim 13\%$ occasionally, confirming the first two constraints.
+Most cases show only a few percent deviations from averaged values, with a maximum deviation of $\sim 16\%$ occasionally, confirming the first two constraints.
 To test the third constraint, we decompose $M$ into symmetric ($M^{\text{s}}=(M+M^{\text{T}})/2$) and antisymmetric ($M^{\text{a}}=(M-M^{\text{T}})/2$) components.
@@
 The experimental data points in (b) were extracted by fitting the corresponding data for $F_{\alpha\gamma}(\phi,t)$ with a sixth-order polynomial.
+The errorbars represent the standard deviations.
 The excellent agreement with experimental data strongly validates the scramblon ansatz.
@@
 For comparison, we examine an alternative mitigation approach from previous literatures that normalizes $F_{\alpha\gamma}(\phi,t)$ by $F(0,t)$ to define $\tilde{F}_{\alpha\gamma}(\phi,t)=F_{\alpha\gamma}(\phi,t)/F_{\alpha\gamma}(0,t)$ (red triangles in Fig.
-4a).
+4a) [19, 15, 9, 48].
 Within the scramblon formalism, this yields:
@@
 Notably, time-reversal-based exponential sensitivity has also been recently exploited to enhance quantum metrology in systems such as cold atoms in cavities [24, 25] and nitrogen vacancy centers [31], where imperfections also suppress this sensitivity at long evolution times.
-Our theory may be extended to such quantum metrology applications, potentially enabling further improvements in sensitivity.
+Our theory may be extended to such quantum metrology applications, potentially enabling further improvements in sensitivity [49].
 Acknowledgements This work is supported by Innovation Program for Quantum Science and Technology 2021ZD0302005 (H.Z.), 2024ZD0300101 (P.Z.), and 2021ZD0303205 (X.P.), National Natural Science Foundation of China No.

### v3

@@
 Therefore, successfully reversing quantum many-body dynamics demands a deep understanding of the underlying structure of quantum information scrambling and chaotic dynamics.
-Here, by using solid-state nuclear magnetic resonance on a macroscopic ensemble of randomly interacting spins, we measure the out-of-time-ordered correlator (OTOC) and validate key predictions of scramblon theory, a universal theoretical framework for information scrambling.
+In this letter, by using solid-state nuclear magnetic resonance on a macroscopic ensemble of randomly interacting spins, we measure the out-of-time-ordered correlator (OTOC) and validate key predictions of scramblon theory, a universal theoretical framework for information scrambling.
 Crucially, this theory enables us to isolate and mitigate errors in the OTOC caused by imperfections in the backward evolution.
@@
 At the same time, the entanglement entropy grows, rendering the system locally a thermal state.
-Intriguingly, one can envision a gedanken experiment: if we apply the reversed Hamiltonian $-\hat{H}$ to this thermal state, is it possible to recover the original low-entanglement state and its local information?
+Intriguingly, one can envision a gedanken experiment: if we apply the reversed Hamiltonian $\hat{H}$ to this thermal state, is it possible to recover the original low-entanglement state and its local information?
 This question echoes a profound analogy in black hole physics: whether information thrown in and thermalized by a black hole can ever be retrieved.
@@
 As illustrated by Fig.
-1a, even infinitesimal imperfections in the engineered reversed Hamiltonian can lead to an exponentially amplified error as evolution time increases, a hallmark of quantum many-body chaos [32, 33, 34, 35, 36].
-Given that no practical implementation can perfectly realize $-\hat{H}$, the inherent chaotic nature of quantum many-body dynamics fundamentally prohibits the retrieval of information from a thermal state.
+1(a), even infinitesimal imperfections in the engineered reversed Hamiltonian can lead to an exponentially amplified error as evolution time increases, a hallmark of quantum many-body chaos [32, 33, 34, 35, 36].
+Given that no practical implementation can perfectly realize $\hat{H}$, the inherent chaotic nature of quantum many-body dynamics fundamentally prohibits the retrieval of information from a thermal state.
 The same reasoning extends analogously to the challenge of recovering information from a black hole.
@@
 (a) Chaotic quantum many-body dynamics: Unitary evolution under $\hat{H}$ scrambles quantum information, transforming a low-entanglement initial state into a highly entangled state.
-Under time-reversed dynamics with perturbed Hamiltonian $-\hat{H}+\delta\hat{H}$, the system exhibits exponential deviation from perfect state recovery.
+Under time-reversed dynamics with perturbed Hamiltonian $\hat{H}+\delta\hat{H}$, the system exhibits exponential deviation from perfect state recovery.
 (b) Microscopic structure of adamantane (C10H16) powder: Each granule contains adamantane molecules arranged in a face-centered cubic lattice.
@@
 (1) into the engineered form Eq.
-(2).
+(2) [37, 8].
 Bottom: Experimental sequence for MQC and OTOC measurements.
@@
 This can be achieved using a well-established protocol for measuring the multiple quantum coherence (MQC) spectrum (Fig.
-1c, bottom panel) [43, 44].
+1(c), bottom panel) [44, 45].
 The measurement involves the reverse evolution of this many-body system, which, of course, is imperfect.
@@
 With this fitting formula, we can extrapolate the measured OTOC to the error-free limit.
-We demonstrate that the OTOC after this error mitigation exhibits the anticipated long-time behavior, in sharp contrast to other methods explored in the previous literature.
+We demonstrate that the OTOC after this error mitigation exhibits the anticipated long-time behavior, in sharp contrast to other methods explored in the previous literature[19, 15, 9].
 Figure 2: Fitting Experimental Data with the Scramblon Ansatz.
@@
 Experimentally, the $95\%$ confidence intervals determined from read-out noise are approximately $10^{-4}$, and therefore the error bars are contained within the data markers.
-The experiment is performed on a powdered adamantane (C10H16) sample, as illustrated in Fig.
-1b and detailed in the supplementary material [45].
+Physical system and experimental protocol—The experiment is performed on a powdered adamantane (C10H16) sample, as illustrated in Fig.
+1(b) and detailed in the supplementary material [46].
 Each adamantane molecule contains 16 nuclear spin-$1/2$ hydrogen nuclei, denoted by $\mathbf{S}_{im}$, where the index $i$ labels the molecules and $m=1,\dots,16$ identifies individual spins within a molecule.
@@
 Then, by applying periodic global radio-frequency pulse sequences to rotating spins, as depicted in Fig.
-1c, a more general effective Hamiltonian can be engineered via the Floquet-Magnus expansion as $\hat{H}_{\text{eff}}=\hat{H}_{0}+\hat{H}_{1}$.
+1(c), a more general effective Hamiltonian can be engineered via the Floquet-Magnus expansion as $\hat{H}_{\text{eff}}=\hat{H}_{0}+\hat{H}_{1}$.
 $\hat{H}_{0}$ represents the leading-order contribution in the expansion, which can be generally written as:
@@
 However, the higher-order terms, denoted by $\hat{H}_{1}$, generally cannot be simultaneously reversed.
-Therefore, without loss of generality, we denote the forward evolution as governed by $\hat{H}=\hat{H}_{0}+\hat{H}_{1}$, and the backward evolution as governed by $-\hat{H}_{0}-\hat{H}_{1}^{\prime}=-\hat{H}-\delta\hat{H}$.
+Therefore, without loss of generality, we denote the forward evolution as governed by $\hat{H}=\hat{H}_{0}+\hat{H}_{1}$, and the backward evolution as governed by $\hat{H}_{0}-\hat{H}_{1}^{\prime}=-\hat{H}-\delta\hat{H}$.
 Here, $\delta\hat{H}$, representing the difference between $\hat{H}_{1}$ and $\hat{H}^{\prime}_{1}$, is the imperfection in the reversed dynamics.
@@
 This is followed by backward evolution as described above.
-Finally, we measure the total magnetization along the $\gamma$-direction $\langle\hat{O}_{\gamma}\rangle$, which equals the correlation function $F_{\alpha\gamma}(\phi,t)$ defined as:
+Finally, we measure the total magnetization along the $\gamma$-direction $\langle\hat{O}_{\gamma}\rangle$, which equals the correlation function $\epsilon F_{\alpha\gamma}(\phi,t)$ defined as:
 $F_{\alpha\gamma}(\phi,t)=\text{Tr}[\hat{O}_{\gamma}(0)\hat{\mathcal{V}}^{\dagger}(t){\rm e}^{-{\rm i}\phi\hat{O}_{\alpha}(t)}\hat{O}_{\gamma}(0){\rm e}^{{\rm i}\phi\hat{O}_{\alpha}(t)}\hat{\mathcal{V}}(t)].$ (3)
-Here, $\hat{\mathcal{V}}(t)\equiv\exp\left(-{\rm i}\mathbf{T}\int_{0}^{t}{\rm d}t^{\prime}\delta\hat{H}(t^{\prime})\right)$ represents the time-evolution operator in the interaction picture, with $\mathbf{T}$ denoting time ordering.
+Here, $\hat{\mathcal{V}}(t)\equiv\mathbf{T}\exp\left(-{\rm i}\int_{0}^{t}{\rm d}t^{\prime}\delta\hat{H}(t^{\prime})\right)$ represents the time-evolution operator in the interaction picture, with $\mathbf{T}$ denoting time ordering.
 All operators $\hat{O}(t)$ and $\delta\hat{H}(t)$ evolve according to the Heisenberg picture under $\hat{H}$.
@@
 The MQC spectrum can be obtained through the Fourier transformation of $F_{\alpha\gamma}(\phi,t)$ with respect to $\phi$ [44, 45].
-We now highlight the relationship between $F_{\alpha\gamma}(\phi,t)$ and OTOCs through two limits: (i) For perfect reversed dynamics ($\delta\hat{H}=0$), $F_{\alpha\gamma}(\phi,t)$ reduces to
+Validating the scramblon theory—We now highlight the relationship between $F_{\alpha\gamma}(\phi,t)$ and OTOCs through two limits: (i) For perfect reversed dynamics ($\delta\hat{H}=0$), $F_{\alpha\gamma}(\phi,t)$ reduces to
 $F_{\alpha\gamma}^{(0)}(\phi,t)=\text{Tr}[\hat{O}_{\gamma}(0){\rm e}^{-{\rm i}\phi\hat{O}_{\alpha}(t)}\hat{O}_{\gamma}(0){\rm e}^{{\rm i}\phi\hat{O}_{\alpha}(t)}],$ (4)
@@
 and this OTOC, also known as the Loschmidt echo, probes imperfections in the reversed dynamics.
-Both cases can be computed by scramblon theory [41], strongly suggesting that the general behavior of $F_{\alpha\gamma}(\phi,t)$, although more challenging to derive, can also be captured by scramblon theory.
+Both cases can be computed by scramblon theory [42, 67], strongly suggesting that the general behavior of $F_{\alpha\gamma}(\phi,t)$, although more challenging to derive, can also be captured by scramblon theory.
 Indeed, as we show in the Supplementary Material [46], an ansatz for $F_{\alpha\gamma}(\phi,t)$ in the small-$\phi$ regime can be derived as:
 $F_{\alpha\gamma}(\phi,t)=\frac{1}{\left(1+a{\rm e}^{\varkappa t}+b\phi^{2}{\rm e}^{\varkappa t}\right)^{2\Delta}},$ (6)
-where $\{a,b,\varkappa,\Delta\}$ are all fitting parameters.
+where $\{a,b,\varkappa,\Delta\}$ are all fitting parameters that depend on $\alpha$ and $\gamma$.
 The $\phi=0$ limit of Eq.
@@
 The three columns correspond to the fitting parameters extracted from three different Hamiltonians: $\hat{H}^{(1)}$, $\hat{H}^{(2)}$, and $\hat{H}^{(3)}$.
-Subfigures a-c in the upper row display the fitting parameters for different values of $\alpha$ and $\gamma$, with green diamonds representing $a\times 50$, yellow squares representing $\Delta$, and blue triangles representing $\varkappa$.
+Subfigures (a)-(c) in the upper row display the fitting parameters for different values of $\alpha$ and $\gamma$, with green diamonds representing $a\times 50$, yellow squares representing $\Delta$, and blue triangles representing $\varkappa$.
 The dashed lines represent $\varkappa$ averaged over $\alpha$ and $\gamma$, and $a$ and $\Delta$ averaged over $\alpha$.
-Subfigures d-f in the lower row show the eigenvalues $\sigma_{1},\sigma_{2},\sigma_{3}$ for the symmetric matrix $M^{\text{s}}$ and the absolute value of eigenvalues $\sigma_{4},\sigma_{5},\sigma_{6}$ for the antisymmetric matrix $M^{\text{a}}$.
+Subfigures (d)-(f) in the lower row show the eigenvalues $\sigma_{1},\sigma_{2},\sigma_{3}$ for the symmetric matrix $M^{\text{s}}$ and the absolute value of eigenvalues $\sigma_{4},\sigma_{5},\sigma_{6}$ for the antisymmetric matrix $M^{\text{a}}$.
 The error bars on the fitting parameters represent the $95\%$ confidence interval, including the standard deviation from each fitting residual $\sigma_{\text{res}}$ and the uncertainty arising from the choice of fitting range $\sigma_{\text{range}}$ [46].
@@
 (6), with three representative fits shown in Fig.
-2 and the complete set of all data available in the Supplementary Material [45].
+2.
+We quantify the goodness of fit using the adjusted $R^{2}$, as detailed in the Supplementary Material [46].
+Discrepancies between experiment and theory may arise from the invalidity of the scramblon theory at very early times $e^{\varkappa t}\lesssim 1$.
 The scramblon theory not only provides the fitting ansatz but also imposes strong constraints on the parameters.
@@
 Eigenvalues shown in panels (d-f) reveal that $M^{\text{s}}$ exhibits only one dominant eigenvalue, and all eigenvalues of $M^{\text{a}}$ are negligible, validating the rank-1 symmetric requirement for $M$.
+It is notable that a strong correspondence between the results for $H^{(1)}_{0}$ and $H^{(2)}_{0}$ is observed.
+This is anticipated, as $H^{(1)}_{0}$ can be transformed into $H^{(2)}_{0}$ via a global rotation ${\rm e}^{-{\rm i}\hat{O}_{z}\pi/4}$.
 Figure 4: Error Mitigation of the Reversed Quantum Many-Body Dynamics.
@@
 The errorbars represent the standard deviations.
-The excellent agreement with experimental data strongly validates the scramblon ansatz.
+Recovering exponential behavior of OTOC and extracting the Lyapunov exponent—The excellent agreement with experimental data strongly validates the scramblon ansatz.
 Within this framework, the parameter $a$ fully captures imperfections in the time-reversed dynamics, as evidenced by $F_{\alpha\gamma}(0,t)\equiv 1$ when $a=0$.
@@
 (6) while keeping other fitted parameters unchanged (blue curve in Fig.
-4a).
+4(a)).
 For comparison, we examine an alternative mitigation approach from previous literatures that normalizes $F_{\alpha\gamma}(\phi,t)$ by $F(0,t)$ to define $\tilde{F}_{\alpha\gamma}(\phi,t)=F_{\alpha\gamma}(\phi,t)/F_{\alpha\gamma}(0,t)$ (red triangles in Fig.
-4a) [19, 15, 9, 48].
+4(a)) [19, 15, 9, 69].
 Within the scramblon formalism, this yields:
 $\tilde{F}_{\alpha\gamma}(\phi,t)=\left(\frac{1+a{\rm e}^{\varkappa t}}{1+a{\rm e}^{\varkappa t}+b\phi^{2}{\rm e}^{\varkappa t}}\right)^{2\Delta}.$ (7)
-Different approaches are compared in Fig. 4a.
+Different approaches are compared in Fig. 4(a).
 To assess the reliability of different approaches, we compute $I_{\alpha\gamma}(\phi)$ from $F_{\alpha\gamma}(\phi,t)$ as:
@@
 Analysis based on the scramblon ansatz reveals: (i) The raw data (Eq.
-(6)) gives $I_{\alpha\gamma}(t)=2b\Delta{\rm e}^{\varkappa t}/(1+a{\rm e}^{\varkappa t})^{1+\Delta}$, which decays at late times due to $\Delta>0$ (green dashed line in Fig.
-4b); (ii) The normalized approach (Eq.
-(7)) yields $I_{\alpha\gamma}(t)=2b\Delta{\rm e}^{\varkappa t}/(1+a{\rm e}^{\varkappa t})$, saturating to $2b\Delta/a$ at long times (red dash-dotted line in Fig.
-4b); and (iii) Only our $a=0$ mitigation produces the physically expected exponential growth $2b\Delta{\rm e}^{\varkappa t}$ (blue solid line in Fig.
-4b), where $\varkappa$ emerges as the quantum Lyapunov exponent.
+(6)) gives $I_{\alpha\gamma}(t)=4b\Delta{\rm e}^{\varkappa t}/(1+a{\rm e}^{\varkappa t})^{1+2\Delta}$, which decays at late times due to $\Delta>0$ (green dashed line in Fig.
+4(b)); (ii) The normalized approach (Eq.
+(7)) yields $I_{\alpha\gamma}(t)=4b\Delta{\rm e}^{\varkappa t}/(1+a{\rm e}^{\varkappa t})$, saturating to $4b\Delta/a$ at long times (red dash-dotted line in Fig.
+4(b)); and (iii) Only our $a=0$ mitigation produces the physically expected exponential growth $4b\Delta{\rm e}^{\varkappa t}$ (blue solid line in Fig.
+4(b)), where $\varkappa$ emerges as the quantum Lyapunov exponent.
+It is also remarkable that only our $a=0$ mitigated result $I_{\alpha\gamma}(t)$ exhibits the symmetric and rank-1 property predicted by the scramblon theory for systems with time-reversal symmetry, because there is a single dominant scramblon mode and its coupling to both $\hat{O}_{\alpha}$ and $\hat{O}_{\gamma}$ operators is symmetric (see Supplementary Materiral [46] for more experimental data on all combinations of $\alpha$ and $\gamma$).
 These distinct behaviors are quantitatively compared with experimental data.
@@
 3 serves as the first exact experimental determination of the Lyapunov exponent in a quantum many-body system with macroscopic degrees of freedom.
-In conclusion, we develop an ansatz based on scramblon theory to describe reversed quantum many-body dynamics in the presence of imperfections.
+Conclusion—In conclusion, we develop an ansatz based on scramblon theory to describe reversed quantum many-body dynamics in the presence of imperfections.
 We validate this ansatz by fitting it to experimental data from an NMR system with macroscopically randomly interacting spins, where the stringent constraints imposed by scramblon theory are convincingly verified by the fitting outputs.
-This approach allows us to mitigate the effects of imperfections and recover error-resilient reversed dynamics.
+These results suggest that the applicability of the scramblon theory extends beyond idealized all-to-all interacting models, remaining valid in realistic systems with microscopic long-range interactions.
+The application of the scramblon theory allows us to mitigate the effects of imperfections and recover error-resilient reversed dynamics.
 Notably, time-reversal-based exponential sensitivity has also been recently exploited to enhance quantum metrology in systems such as cold atoms in cavities [24, 25] and nitrogen vacancy centers [31], where imperfections also suppress this sensitivity at long evolution times.
-Our theory may be extended to such quantum metrology applications, potentially enabling further improvements in sensitivity [49].
-Acknowledgements This work is supported by Innovation Program for Quantum Science and Technology 2021ZD0302005 (H.Z.), 2024ZD0300101 (P.Z.), and 2021ZD0303205 (X.P.), National Natural Science Foundation of China No.
+Our theory may be extended to such quantum metrology applications, potentially enabling further improvements in sensitivity.
+Acknowledgements—This work is supported by Quantum Science and Technology-National Science and Technology Major Project 2021ZD0302005 (H.Z.), 2024ZD0300101 (P.Z.), and 2021ZD0303205 (X.P.), National Natural Science Foundation of China No.
 12488301 (H.Z.), No.
@@
 We thank experimental assistance from the Instruments Center for Physical Science, University of Science and Technology of China.
-Data and Code Availability The raw NMR experimental data and all codes are available at [50].
+Data Availability—The data that support the findings of this study are available at Zenodo (https://doi.org/10.5281/zenodo.16142455) [70].
+Further data are available from the corresponding authors upon reasonable request.